OpenStudy (anonymous):

a room has a perimeter of 48 ft. The width is 2/3 of the length. What are the dimensions of the room?

OpenStudy (amistre64):

sounds like the garage

OpenStudy (anonymous):

Perimeter= 2l+2w So....48=2l+2(2/3)l 48=(8/3)l l=18

OpenStudy (anonymous):

w=5

OpenStudy (anonymous):

oops i mean....w=6

OpenStudy (anonymous):

how did you get l=18